A factory worker pushes a 28.1 kg crate a distance of 4.1 m along a level floor at constant velocity by pushing downward at an angle of 31 ∘ below the horizontal. The coefficient of kinetic friction between the crate and floor is 0.26.

Part A: What magnitude of force must the worker apply to move the crate at constant velocity?

Part B: How much work is done on the crate by this force when the crate is pushed a distance of 4.1 m ?

Part C: How much work is done on the crate by friction during this displacement?

Part D: How much work is done by the normal force?

Part E: How much work is done by gravity?
Answer #1

A) if he pushes with constant velocity ,Net force will be zero

Force in vertical direction = F sin 31

mg + F sin 31 = N

N = 28.1*9.8 + 0.515F

N = 275.38 + 0.515F

friction force = horizontal force

uN = F cos 31

0.26*(275.38 + 0.515F) = F cos 31

F = 99 N

B) work done by the force =F*S*cos 31 = 99*4.1*cos 31 = 347.9 J

C) Work done by friction = F*S*cos 180

= (99*cos 31)*4.1 * cos 180 = -347.9 J

D) work done by normal force = F*S *cos 90 = 0

E) work done by gravity = F*S*cos 90 = 0